Mr LEO’s father has made a will to give him an amount of $500,000 in 10 years .So Leo wants to borrow an amount from the bank to set up a fashion boutique in the nearest town. The business planner told him that he does not expect the present discount rate to change in these 10 years. The following information is given regarding the discount rate

  GDP growth of the country

1.2%

Beta

1.8

Index in January 2020

16,570

Index in December 2020

17,590

Inflation

2.5%

 REQUIRED

How much should he borrow from the bank now if he is planning to pay back the loan with the money he will get in 10 years of business return?


 



1
Expert's answer
2021-06-28T18:27:04-0400

step 1

The question is based on the concept of calculating the value of a loan in terms of present value according to repay from the future value of the current investment.

Step 2

Expected Inflow at end of 10 year= $ 500,000.

return from index(equity) = "\frac{17590-16570}{16570}" =6.156%

risk free rate = (1+1.2%)"\ast" (1+2.5%)-1 = 3.73%

discount rate = risk free+ "\beta" "\ast" (equity premium)

discount rate = 3.73% + 1.8"\ast" (6.156%-3.73%)=8.096%





Current Loan amount : 



loan ammount = "\frac{NFV}{(1+discount rate)10}=\frac{500000}{(1+8.096)10}" =$229,542.496

Current laon = $ 229,542.50